# LLM-Enhanced User-Item Interactions: Leveraging Edge Information for Optimized Recommendations

### v1

@@ 6. Conclusion
We tackle a key issue in recommendation systems: how to integrate LLM and graph structures into recommendations. To this end, we propose a graph attentive LLM generative recommender system. By introducing new prompting methods and graph structured attention mechanisms, we can effectively integrate the complex relationships and background information between users and items into the model. We first design a natural language prompt that can reflect the relationship between users and items and embedded the 2-order relationship between items into it. Next, we improved the attention mechanism of LLM to model complex graph structure information. Through experiments, we validate the effectiveness of our method. The experimental results show that our model has significantly improved recommendation accuracy and personalization compared to traditional recommendation systems. Considering these innovations, our approach provides a new technological path for developing more efficient and intelligent recommendation systems. Meanwhile, these methods demonstrate new perspectives and ideas in applying LLM to recommendation systems and a wider range of fields. This promotes the development of recommendation systems and provides strong support and inspiration for using LLM in various complex application scenarios.

### v2

@@ 6. Conclusion
We tackle a key issue in recommender systems: how to integrate LLM and graph structures into recommendations. To this end, we propose a graph attentive LLM generative recommender system. By introducing new prompting methods and graph-structured attention mechanisms, we can effectively integrate the complex relationships and background information between users and items into the model. We first designed a natural language prompt that can reflect the relationship between users and items and embed the 2-order relationship between items into it. Next, we improved the attention mechanism of LLM to model complex graph structure information. Through experiments, we validate the effectiveness of our method. The experimental results show that our model has significantly improved recommendation accuracy and personalization compared to traditional recommender systems. Considering these innovations, our approach provides a new technological path for developing more efficient and intelligent recommender systems. Meanwhile, these methods demonstrate new perspectives and ideas in applying LLM to recommender systems and a wider range of fields. This promotes the development of recommender systems and provides strong support and inspiration for using LLM in various complex application scenarios.